# Estimation of Optimal Causal Bounds via Covariate-Assisted Optimal Transport

### v1

@@ 5.2 Real Data
In this section, we evaluate our method against existing approaches using real-world data from the Student Achievement and Retention (STAR) Demonstration Project [angrist2009incentives], an initiative designed to assess the impact of scholarship incentive programs on academic performance. In the STAR study, the treatment is access to a scholarship incentive and was randomly assigned. Academic performance is measured by the GPA recorded at the end of the first academic year. In addition to the treatment indicator and outcome variable, the dataset includes baseline GPA (measured prior to treatment assignment), which serves as a key covariate due to its strong predictive influence on academic outcomes.
@@
$\rho:=\frac{\mathbb{E}[Y_{i}(1)Y_{i}(0)]-\mathbb{E}[Y_{i}(1)]\mathbb{E}[Y_{i}(0)]}{\sqrt{\mathrm{Var}(Y_{i}(0))\mathrm{Var}(Y_{i}(1))}}.$
The unidentifiable part of $\rho$ arises from the estimand $\mathbb{E}[Y_{i}(1)Y_{i}(0)]$. Then, to obtain the corresponding $V_{\textup{c}}$ value for $\rho$, we could utilize the cost function $h(y_{0},y_{1})=(y_{0}+y_{1})^{2}$.
To ensure the PI bounds are accessible at the population level, we generate hypothetical data from a model fitted to the real data. Particularly, we fit the real data using the model
@@
The results are collected in Table 1. We observe that the adapted COT value estimator consistently outperforms the other methods. The ridge regression–based DualBounds method also performs competitively, likely due to the outcome–covariate relationship being approximately linear in the underlying data-generating process.
Table 1: Relative error comparison across methods and sample sizes. Here, $N$ denotes the total number of units, with 50% assigned to the treatment group. Each reported error is the average over 500 Monte Carlo simulation runs, and the value in parentheses indicates the associated standard error. For the row labels, “adapt” refers to the adapted COT value estimator, “ridge” refers to the ridge regression–based DualBounds method, and “knn” refers to the KNN–based DualBounds method.
Method $N=400$ $N=600$ $N=1200$ $N=1600$
ridge $4.0\text{e}{-2}$ (1.3e-3) $3.2\text{e}{-2}$ (1.0e-3) $2.9\text{e}{-2}$ (0.9e-3) $2.7\text{e}{-2}$ (0.8e-3)
knn $5.1\text{e}{-2}$ (1.6e-3) $4.5\text{e}{-2}$ (1.3e-3) $4.5\text{e}{-2}$ (1.1e-3) $4.4\text{e}{-2}$ (1.0e-3)
adapt $\mathbf{3.6e\text{-}2~{}(1.2e\text{-}3)}$ $\mathbf{2.9e\text{-}2~{}(1.0e\text{-}3)}$ $\mathbf{2.6e\text{-}2~{}(0.8e\text{-}3)}$ $\mathbf{2.5e\text{-}2~{}(0.8e\text{-}3)}$
6 Conclusion and Discussion
In this paper, we obtain the optimal covariate-assisted partial identification sets for a class of causal estimands by solving COT problems. Our finite-sample estimator of the COT value avoids nuisance function estimation, does not require well-specified models, and enjoys a provable statistical convergence rate. Furthermore, our adapted COT value estimator answers the question proposed in the discussion of [lin2025tightening] on how to combine the adapted Wasserstein distance into the COT value estimation. Specifically, our adapted estimator is their $V_{\text{causal}}(\eta)$ when $\eta$ approached infinity.
A direction for future research is to replace the current equal-sized cells with data-driven counterparts. Specifically, we can apply tree-based methods to the control and treatment group to identify regions where the distribution of $Y_{i}(0)$ is approximately homogeneous, and then project the covariates of control units onto the centers of these regions. Such data-driven partitions avoid splitting on covariates that are not influential to the outcome distribution and thus mitigate the curse of dimensionality. Another future direction is to move beyond estimation and conduct inference for the PI interval. One possible approach is deriving the asymptotic distribution of our estimator, potentially leveraging recent central limit theorem results for optimal transport mappings [manole2023central], at least for moderately large $d_{Y}$, $d_{Z}$.
Acknowledgements
The material in this paper is partly supported by the Air Force Office of Scientific Research under award number FA9550-20-1-0397 and ONR N000142412672. Support from NSF 2229012, 2312204, 2403007 is also gratefully acknowledged.
[…]
Appendix
Organization. The Appendix is organized as follows. Section A reviews technical background knowledge. Section B provides proofs for the results in Section 2. Section C provides proofs for the results in Section 3. Section D provides proofs for the results in Section 4. Section E collects proofs for supporting lemmas.
Appendix A Technical Background
A.1 Regular Kernel
@@
Definition A.1 (Regular kernel).
@@
Proposition A.1 (Disintegration [kallenberg1997foundations, Theorem 5.3, 5.4]).
@@
A.2 Wasserstein vs. Adapted Wasserstein: A Distance Metric Comparison
@@
Definition A.2 (Wasserstein distance).
@@
Lemma A.1 (Triangle inequality).
@@
Definition A.3 (Adapted Wasserstein distance).
@@
$\mathds{W}_{\textup{bc}}(\mu,\nu)\triangleq\min_{\pi\in\Pi_{\textup{bc}}(\mu,\nu)}\mathbb{E}_{\pi}[\left\|{X(0)-X(1)}\right\|_{{2}}],$
@@
$\Pi_{\textup{bc}}(\mu,\nu)\triangleq$$\left\{\pi\in\Pi(\mu,\nu):\text{under}~{}\pi,~{}Z(1)\perp\!\!\!\perp Y(0)\text{ conditioned on }Z(0),\right.$
@@
The following proposition is stated to maintain consistency with Definition 2.1 introduced earlier.
Proposition A.2 (Reformulation of adapted Wasserstein distance; see, e.g. [backhoff2017causal, Proposition 5.2]).
@@
$\mathds{W}_{\textup{bc}}(\mu,\nu)=\min_{\pi\in\Pi(\mu_{Z},\nu_{Z})}\int\left\|{z-z^{\prime}}\right\|_{{2}}+\mathds{W}_{1}(\mu_{Y}^{z},\nu_{Y}^{z^{\prime}})~{}{\rm d}\pi(z,z^{\prime}).$
Since $\Pi_{\textup{bc}}(\mu,\nu)\subseteq\Pi(\mu,\nu)$, clearly $\mathds{W}_{1}\leq\mathds{W}_{\textup{bc}}$, but the reverse may not hold. Consider an i.i.d. sample $(X_{i}=(Z_{i},Y_{i}),i\in[n])$ drawn from a distribution $\mu$, and let $\hat{\mu}_{n}\triangleq\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}$. It is well known that $\mathds{W}_{1}(\hat{\mu}_{n},\mu)$ converges to zero almost surely as $n$ goes to infinity. However, as shown in [pflug2016empirical, Proposition 1], $\mathds{W}_{\textup{bc}}(\hat{\mu}_{n},\mu)$ may fail to converge to zero, and a remedy to ensure convergence is to replace $\hat{\mu}_{n}$ with its convolution with a suitable kernel function. The intuition behind this approach is to ensure that the kernel of the convoluted empirical distribution becomes smooth in the variable it is conditioned on. This idea aligns with the findings of [blanchet2024bounding], which show that for distributions with Lipschitz kernels, the weak topology and the topology induced by the adapted Wasserstein distance are equivalent.
However, since the convolved empirical distribution is not discrete, its practical implementation requires an additional sampling step for computation. To address this, [backhoff2022estimating] propose a discrete alternative—the adapted empirical distribution—which mitigates the convergence issue by assigning the sample $(X_{i})_{i\in[n]}$ to a finite number of cells, which is the approach we adopt in this paper.
A.3 Founier and Guillin’s Result
This section provides the convergence rate of the empirical distribution under the Wasserstein distance, as established by [fournier2015rate].
Suppose that the sample $(X_{i},i\in[n])$ i.i.d. follows $\mu$ and denote $\hat{\mu}_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}$. Theorem 1 in [fournier2015rate] characterizes a nearly sharp convergence rate for empirical Wasserstein distances.
Theorem A.1 (Empirical Wasserstein convergence rate, Theorem 1 [fournier2015rate]).
@@ A.3 Founier and Guillin’s Result
Corollary A.1 ($k>2$).
Under the same conditions of Theorem A.1, if $k>2$, then
@@
Appendix B Proofs of Section 2
B.1 Proof of Proposition 2.1
Proof of Proposition 2.1.
@@ Appendix B Proofs of Section 2
where we use Assumption 2.2 in the second last equality. ∎
B.2 Proof of Proposition 2.2
Remark B.1.
In this proof, we will use the notions of Borel measurability, lower semianalyticity, and universal measurability, which are discussed in detail in [bertsekas1978stochastic, Section 7.7].
Proof of Proposition 2.2.
@@
Note that the maps $z\mapsto\mu_{Y(0)}^{z}$ and $z\mapsto\mu_{Y(1)}^{z}$ are both Borel since $\mu_{Y(0)}^{z},\mu_{Y(1)}^{z}$ are regular kernels (see Section A.1). Further, $\{(p,q,\gamma):\gamma\in\Pi(p,q)\}$ is weakly closed. Then, the set
@@
As a result, $\mathds{W}_{h}(\mu_{Y(0)}^{z},\mu_{Y(1)}^{z})$ is the infimum of the function $\mathcal{L}_{h}$ over the fiber of $D$ at $z$. By [bertsekas1978stochastic, Proposition 7.47], the map $z\mapsto\mathds{W}_{h}(\mu_{Y(0)}^{z},\mu_{Y(1)}^{z})$ is l.s.a. and in particular universally measurable. Therefore, the RHS $\int\mathds{W}_{h}(\mu_{Y(0)}^{z},\mu_{Y(1)}^{z}){\rm d}\mu_{Z}(z)+\epsilon.$ is well-defined.
@@
LHS $\leq$ RHS: For any $\epsilon>0$ By [bertsekas1978stochastic, Proposition 7.50(b)], there is a universally measurable $\epsilon$-optimizer for
@@
Then, apply [bertsekas1978stochastic, Proposition 7.45], we can construct a measure $\pi_{\epsilon}\in\mathcal{P}(\mathcal{Y}^{2}\times\mathcal{Z})$ such that
@@
Appendix C Proofs of Section 3
C.1 Proof of Theorem 3.1
Lemma C.1 ($\Pi_{\textup{c}}$ is compact).
@@ Appendix C Proofs of Section 3
For (ii), by definition of $V_{\textup{c}}$, for any $\pi\in\Pi_{\textup{c}}$, $\mathbb{E}_{\pi}[h(Y(0),Y(1))]\geq V_{\textup{c}}$, indicating that $V_{\textup{c}}$ is a valid lower bound for $V^{*}$. Similarly, $\tilde{V}_{\textup{c}}$ is a valid upper bound for $V^{*}$.
Further, since $h$ is bounded and lower semicontinuous, $\mathbb{E}_{\pi}[h(Y(0),Y(1))]$ is a lower semicontinuous function in $\pi$ with respect to the weak topology ([villani2009optimal, Lemma 4.3]). By Lemma C.1, $\Pi_{\textup{c}}$ is compact, thus the minimum $V_{\textup{c}}$ can be attained. Similarly, the maximum $\tilde{V}_{\textup{c}}$ can also be attained since $h$ is upper semicontinuous. Let $\pi_{\textup{c}}$ (resp. $\tilde{\pi}_{\textup{c}}$) be the optimal coupling corresponding to $V_{\textup{c}}$ (resp. $\tilde{V}_{\textup{c}}$).
Finally, for any $v=(1-\theta)V_{\textup{c}}+\theta\tilde{V}_{\textup{c}}$, where $\theta\in[0,1]$, there is a coupling $\pi_{v}=(1-\theta)\pi_{\textup{c}}+\theta\tilde{\pi}_{\textup{c}}$ such that $\pi_{v}\in\Pi_{\textup{c}}$ (since $\Pi_{\textup{c}}$ is a convex set) and $v=\mathbb{E}_{\pi_{v}}[h(Y(0),Y(1))]$. As a result, $[V_{\textup{c}},\tilde{V}_{\textup{c}}]$ is the tightest valid interval for possible values of $\mathbb{E}_{\pi}[h(Y(0),Y(1))]$, where $\pi$ is aligned with $\mu_{Y(0),Z},\mu_{Y(1),Z}$. ∎
Remark C.1 (Unbounded $h$).
The results of Theorem 3.1 can be extended to any continuous function $h$ such that, there are functions $a(y)\in L^{1}(\mu_{Y(0)}),b(y)\in L^{1}(\mu_{Y(1)})$ and $|h(y_{0},y_{1})|\leq a(y_{0})+b(y_{1})$. For the proof, we can consider $\tilde{h}=h(y_{0},y_{1})-(a(y_{0})+b(y_{1}))$ for the lower bound and $\tilde{h}^{\prime}=a(y_{0})+b(y_{1})-h(y_{0},y_{1})$ for the upper bound.
C.2 Proof of Theorem 3.2
Lemma C.2 (Lipschitz continuity of $\mathds{W}_{h}$).
Under Assp. 3.3, $\mathds{W}_{h}$ is $L_{h}$-Lipschitz with respect to the 1-Wasserstein distance, i.e., for any $(\mu,\nu,\mu^{\prime},\nu^{\prime})\in\mathcal{P}$,
$\left|\mathds{W}_{h}(\mu,\nu)-\mathds{W}_{h}(\mu^{\prime},\nu^{\prime})\right|\leq L_{h}\left(\mathds{W}_{1}(\mu,\mu^{\prime})+\mathds{W}_{1}(\nu,\nu^{\prime})\right).$
Proof of Theorem 3.2.

### v2

@@ Appendix A More Details on the Real Data Experiment
In this section, we present more details of the real data experiment in Section 5.2, which is based on the Student Achievement and Retention (STAR) Demonstration Project [1], an initiative designed to assess the impact of scholarship incentive programs on academic performance. Experiments were run on a MacBook Air (Apple M3, 8-core CPU, 16 GB RAM), without GPU acceleration.
In the STAR study, the treatment is access to a scholarship incentive and was randomly assigned. Academic performance is measured by the GPA recorded at the end of the first academic year. In addition to the treatment indicator and outcome variable, the dataset includes baseline GPA (measured prior to treatment assignment), which serves as a key covariate due to its strong predictive influence on academic outcomes.
@@
$\rho\triangleq\frac{\mathbb{E}[Y_{i}(1)Y_{i}(0)]-\mathbb{E}[Y_{i}(1)]\mathbb{E}[Y_{i}(0)]}{\sqrt{\mathrm{Var}(Y_{i}(0))\mathrm{Var}(Y_{i}(1))}}.$
The unidentifiable part of $\rho$ arises from the estimand $\mathbb{E}[Y_{i}(1)Y_{i}(0)]$. Then, to obtain the corresponding $V_{\textup{c}}$ value for $\rho$, we could utilize the cost function $h(y_{0},y_{1})=(y_{0}+y_{1})^{2}$. To ensure the PI bounds are accessible at the population level, we generate hypothetical data from a model fitted to the real data. Particularly, we fit the real data using the model $Y(w)=f_{w}(Z(w))+\mathcal{N}(0,\sigma_{w}^{2})\ w=0,1,$ and subsequently generate data based on the estimated functions $f_{w}$ and noise level $\sigma_{w}$.
To ensure the true (population-level) PI bounds are accessible, we generate hypothetical data from a model fitted to the real data. Particularly, we fit the real data using the model
@@
Appendix B Supplementary Experiments
B.1 Robustness to Covariate Distributional Shift
In this section, we consider a setting in which the covariates of the treatment group are shifted. Let $\mu_{Z(0)},\mu_{Z(1)}$ denote the (marginal) generating distribution of $Z(0),Z(1)$. Specifically, we set $\mu_{Z}=\mu_{Z(0)}=\mathcal{N}(0,1)$ and perturb the treatment covariate distribution as $\mu_{Z(1)}=\mathcal{N}(0,1)+\eta\varepsilon$, where $\varepsilon$ is an independent $\mathcal{N}(0,1)$ noise and $\eta$ controls the magnitude of the perturbation. Using the same three models described in Section 5.2, we evaluate the estimation accuracy of the adapted COT value estimator under varying levels of $\eta$. The results are reported in Figure 4. We observe that the adapted COT value estimator remains stable across small perturbation levels ($\eta\in[0,0.1]$), illustrating its robustness to covariate distribution shift (see also Theorem 5). Intuitively, this robustness stems from the projection of covariates onto a finite set of representative values: small perturbations in the covariates are often absorbed during coarsening, as they rarely alter cell assignments. Moreover, the use of optimal coupling to match the discretized marginals helps mitigate the impact of systematic perturbations such as location shifts.
B.2 Comparison with a Fréchet–Hoeffding-Type Approach
When the outcomes are one-dimensional, the copula models [39], and the Fréchet–Hoeffding bounds [25, 37, 20, 31] can provide partial identification bounds in closed forms without solving an OT-type optimization problem. Particularly, for the case $d_{Y}=1$ and $h(y_{0},y_{1})=(y_{0}-y_{1})^{2}$, Hoeffding-type results yield a closed-form optimizer for (4). (Note that Hoeffding-type bounds do not provide consistent partial identification sets for $d_{Y}>1$, whereas our procedure does.)
Proposition 3 (Fréchet–Hoeffding bounds for $h(y_{0},y_{1})=(y_{0}-y_{1})^{2}$).
Let $F_{w}(\cdot|z)\ w=0,1$ be the marginal CDFs of $Y_{w}$ conditioned on $Z=z$ on $\mathbb{R}$ with finite second moments, and denote their quantile functions by $F_{w}^{-1}(u|z)=\inf\{x:F_{w}(x|z)\geq u\}$ for $w=0,1$. Then, for any coupling $\pi$ of $(Y_{0},Y_{1},Z)$ with these marginals,
$\int_{\mathcal{Z}}\int_{0}^{1}\!\big(F_{0}^{-1}(u|z)-F_{1}^{-1}(u|z)\big)^{2}\,{\rm d}u{\rm d}z$ $\;\leq\;\mathbb{E}_{\pi}\!\big[(Y_{0}-Y_{1})^{2}\big]\;\leq\;\int_{\mathcal{Z}}\int_{0}^{1}\!\big(F_{0}^{-1}(u|z)-F_{1}^{-1}(1-u|z)\big)^{2}\,{\rm d}u{\rm d}z.$
Conditioned on $Z=z$, the lower bound is attained by the comonotone coupling $Y_{0}=F_{0}^{-1}(U|z)$, $Y_{1}=F_{1}^{-1}(U|z)$, and the upper bound by the countermonotone coupling $Y_{0}=F_{0}^{-1}(U|z)$, $Y_{1}=F_{1}^{-1}(1-U|z)$, where $U\sim\mathrm{Uniform}(0,1)$.
Based on the above result, a Fréchet-Hoeffding-Type approach [2, 6] for estimating $V_{\textup{c}}$ proceeds as follows:
  1. (i)
    Quantile regression: Within the control and treatment groups, estimate the conditional quantile functions by running quantile regressions. For example, in the following, we use a non-parametric quantile regression method, random forest quantile regression (python package: quantile_forest [29] [^6] https://pypi.org/project/quantile-forest/).
  2. (ii)
    Fréchet–Hoeffding bound: For each covariate value (in either the control or treatment group), use the estimated conditional quantiles to compute the squared loss.
  3. (iii)
    Aggregation: Approximate the overall value by taking the sample average of the values of the conditional squared loss derived in (ii) over the distinct covariate values.
To compare our adapted COT value estimator and the above Fréchet–Hoeffding–based approach, we consider the scale model (Section 5.2 (c)) with the same model parameter as there. We run a Monte Carlo simulation with 200 repetitions. The mean absolute error (standard error) results are collected in Table 3 (In this setting we could compute the oracle true value (=1.92) of the PI bounds with high accuracy). The results show that our method is comparable with the Fréchet–Hoeffding–based method in this special one-dimensional outcome setting.
Table 3: Comparison of our method and Fréchet–Hoeffding–based method using the scale model.
N = 100 N = 500 N = 1000 N = 1500
Our Method 0.356 (0.025) 0.176 (0.011) 0.138 (0.007) 0.122 (0.007)
F–H Based 0.276 (0.019) 0.180 (0.009) 0.156 (0.007) 0.148 (0.006)
Appendix C Technical Background
C.1 Regular Kernel
@@
Definition 5 (Regular kernel).
@@
Proposition 4 (Disintegration [30, Theorem 5.3, 5.4]).
@@
C.2 Wasserstein vs. Adapted Wasserstein: A Distance Metric Comparison
@@
Definition 6 (Wasserstein distance).
@@
Lemma 1 (Triangle inequality).
@@
Definition 7 (Adapted Wasserstein distance).
@@
$\mathds{W}_{\textup{a}}(\mu,\nu)\triangleq\min_{\pi\in\Pi_{\textup{a}}(\mu,\nu)}\mathbb{E}_{\pi}[\left\|{X(0)-X(1)}\right\|_{{2}}],$
@@
$\Pi_{\textup{a}}(\mu,\nu)\triangleq$ $\left\{\pi\in\Pi(\mu,\nu):\text{under}~\pi,~Z(1)\perp\!\!\!\perp Y(0)\text{ conditioned on }Z(0),\right$.
@@
The following proposition is stated to maintain consistency with Definition 1 introduced earlier.
Proposition 5 (Reformulation of adapted Wasserstein distance; see, e.g. [4, Proposition 5.2]).
@@
$\mathds{W}_{\textup{a}}(\mu,\nu)=\min_{\pi\in\Pi(\mu_{Z},\nu_{Z})}\int\left\|{z-z^{\prime}}\right\|_{{2}}+\mathds{W}_{1}(\mu_{Y}^{z},\nu_{Y}^{z^{\prime}})~{\rm d}\pi(z,z^{\prime}).$
Since $\Pi_{\textup{a}}(\mu,\nu)\subseteq\Pi(\mu,\nu)$, clearly $\mathds{W}_{1}\leq\mathds{W}_{\textup{a}}$, but the reverse may not hold. Consider an i.i.d. sample $(X_{i}=(Z_{i},Y_{i}),i\in[n])$ drawn from a distribution $\mu$, and let $\hat{\mu}_{n}\triangleq\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}$. It is well known that $\mathds{W}_{1}(\hat{\mu}_{n},\mu)$ converges to zero almost surely as $n$ goes to infinity. However, as shown in [41, Proposition 1], $\mathds{W}_{\textup{a}}(\hat{\mu}_{n},\mu)$ may fail to converge to zero, and a remedy to ensure convergence is to replace $\hat{\mu}_{n}$ with its convolution with a suitable kernel function. The intuition behind this approach is to ensure that the kernel of the convoluted empirical distribution becomes smooth in the variable it is conditioned on. This idea aligns with the findings of [11], which show that for distributions with Lipschitz kernels, the weak topology and the topology induced by the adapted Wasserstein distance are equivalent.
However, since the convolved empirical distribution is not discrete, its practical implementation requires an additional sampling step for computation. To address this, [3] propose a discrete alternative—the adapted empirical distribution—which mitigates the convergence issue by assigning the sample $(X_{i})_{i\in[n]}$ to a finite number of cells, which is the approach we adopt in this paper.
C.3 Fournier and Guillin’s Result
This section provides the convergence rate of the empirical distribution under the Wasserstein distance, as established by [23].
Suppose that the sample $(X_{i},i\in[n])$ i.i.d. follows $\mu$ and denote $\hat{\mu}_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}$. Theorem 1 in [23] characterizes a nearly sharp convergence rate for empirical Wasserstein distances.
Theorem 7 (Empirical Wasserstein convergence rate, Theorem 1 [23]).
@@ C.3 Fournier and Guillin’s Result
Corollary 1 ($k>2$).
Under the same conditions of Theorem 7, if $k>2$, then
@@
Appendix D Proofs of Section 2
D.1 Proof of Proposition 1
Proof of Proposition 1.
@@ D.1 Proof of Proposition 1
where we use Assumption 2 in the second last equality. ∎
D.2 Proof of Proposition 2
Remark 1.
In this proof, we will use the notions of Borel measurability, lower semianalyticity, and universal measurability, which are discussed in detail in [9, Section 7.7].
Proof of Proposition 2.
@@ D.2 Proof of Proposition 2
Note that the maps $z\mapsto\mu_{Y(0)}^{z}$ and $z\mapsto\mu_{Y(1)}^{z}$ are both Borel since $\mu_{Y(0)}^{z},\mu_{Y(1)}^{z}$ are regular kernels (see Section C.1). Further, $\{(p,q,\gamma):\gamma\in\Pi(p,q)\}$ is weakly closed. Then, the set
@@
As a result, $\mathds{W}_{h}(\mu_{Y(0)}^{z},\mu_{Y(1)}^{z})$ is the infimum of the function $\mathcal{L}_{h}$ over the fiber of $D$ at $z$. By [9, Proposition 7.47], the map $z\mapsto\mathds{W}_{h}(\mu_{Y(0)}^{z},\mu_{Y(1)}^{z})$ is l.s.a. and in particular universally measurable. Therefore, the RHS $\int\mathds{W}_{h}(\mu_{Y(0)}^{z},\mu_{Y(1)}^{z}){\rm d}\mu_{Z}(z)+\epsilon.$ is well-defined.
@@
LHS $\leq$ RHS: For any $\epsilon>0$ By [9, Proposition 7.50(b)], there is a universally measurable $\epsilon$-optimizer for
@@
Then, apply [9, Proposition 7.45], we can construct a measure $\pi_{\epsilon}\in\mathcal{P}(\mathcal{Y}^{2}\times\mathcal{Z})$ such that
@@
Appendix E Proofs of Section 3
E.1 Proof of Theorem 1
Lemma 3 ($\Pi_{\textup{c}}$ is compact).
@@ E.1 Proof of Theorem 1
For (ii), by definition of $V_{\textup{c}}$, for any $\pi\in\Pi_{\textup{c}}$, $\mathbb{E}_{\pi}[h(Y(0),Y(1))]\geq V_{\textup{c}}$, indicating that $V_{\textup{c}}$ is a valid lower bound for $V^{*}$. Further, since $h$ is bounded and lower semicontinuous, $\mathbb{E}_{\pi}[h(Y(0),Y(1))]$ is a lower semicontinuous function in $\pi$ with respect to the weak topology ([49, Lemma 4.3]). By Lemma 3, $\Pi_{\textup{c}}$ is compact, thus the minimum $V_{\textup{c}}$ can be attained.
For (iii), for any $v=(1-\theta)V_{\textup{c}}+\theta\tilde{V}_{\textup{c}}$, where $\theta\in[0,1]$, there is a coupling $\pi_{v}=(1-\theta)\pi_{\textup{c}}+\theta\tilde{\pi}_{\textup{c}}$ such that $\pi_{v}\in\Pi_{\textup{c}}$ (since $\Pi_{\textup{c}}$ is a convex set) and $v=\mathbb{E}_{\pi_{v}}[h(Y(0),Y(1))]$. Then, the PI set is convex and thus is an interval.
Since now $h$ is continuous, then by (ii), $\tilde{V}_{\textup{c}}$ is the tightest upper bound for the PI set. As a result, the PI set is equal to $V_{\textup{c}},\tilde{V}_{\textup{c}}]$. ∎
Remark 2 (Unbounded $h$).
The results of Theorem 1 can be extended to possibly unbounded function $h$ such that, there are functions $a(y)\in L^{1}(\mu_{Y(0)}),b(y)\in L^{1}(\mu_{Y(1)})$ ($f\in L^{1}(\mu)$ denotes that $\int|f|{\rm d}\mu<\infty$) and $|h(y_{0},y_{1})|\leq a(y_{0})+b(y_{1})$. For the proof, we can consider $\tilde{h}=h(y_{0},y_{1})-(a(y_{0})+b(y_{1}))$ for the lower bound and $\tilde{h}^{\prime}=a(y_{0})+b(y_{1})-h(y_{0},y_{1})$ for the upper bound.
E.2 Proof of Theorem 2
Lemma 4 (Uniformly continuous kernel).
Suppose that $\mathcal{Y},\mathcal{Z}$ are both compact and $g_{w}(z)=\mu_{Y(w)}^{z}:\mathcal{Z}\rightarrow\mathcal{P}(\mathbb{R}^{d_{Y}})\ w=0,1$ are continuous under the weak topology. Then, for any $\delta>0$, there is a constant $C(\delta)$ such that $\mathds{W}_{1}\left(\mu_{Y(k)}^{z_{0}},\mu_{Y(k)}^{z_{1}}\right)\leq\delta+C(\delta)\|z_{0}-z_{1}\|_{2}$.
Lemma 5 (Lipschitz continuity of $\mathds{W}_{h}$).
For any $\epsilon\geq 0$, suppose that $h:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$ satisfies that for any $y,y^{\prime}\in\mathcal{Y}$,
$\left|h(y)-h(y^{\prime})\right|\leq L_{h}\left\|{y-y^{\prime}}\right\|_{{2}}+\epsilon.$ (10)
Then, for any $(\mu,\nu,\mu^{\prime},\nu^{\prime})\in\mathcal{P}(\mathcal{Y})$,
$\left|\mathds{W}_{h}(\mu,\nu)-\mathds{W}_{h}(\mu^{\prime},\nu^{\prime})\right|\leq L_{h}\left(\mathds{W}_{1}(\mu,\mu^{\prime})+\mathds{W}_{1}(\nu,\nu^{\prime})\right)+2\epsilon.$
Proof of Theorem 2.